How would I solve the equation 2t^2+8t-64=0 using the Factoring Method?
I've gotten the problem down to 2(t-4)(x+8) but then sort of crashed...

Answer by jim_thompson5910(35256)

If you meant to say 2(t-4)(t+8), then you have the correct factorization.

So 2(t-4)(t+8) = 0, which means that t-4=0 or t+8=0 after we use the zero product property.

So t-4=0 ---> t = 4 or t+8=0 ---> t=-8


This means that the solutions are t = 4 or t = -8
--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------